OpenStudy (anonymous):

how do you "evaluate each expression" 81^1/4

OpenStudy (anonymous):



OpenStudy (anonymous):

\[81^\frac{ 1 }{ 4 } = \sqrt[4]{81}\]

OpenStudy (anonymous):

can you please explain what it means to "evaluate" that expression? thanks

OpenStudy (mathstudent55):

Use this rule: \( a ^ {\frac{1}{n} = \sqrt[n]{a}} \)

OpenStudy (mathstudent55):

Evaluate simply means to calculate. For example evaluate \(\sqrt{9} \) simply means to calculate \(\sqrt{9} \) which is 3.

OpenStudy (anonymous):

ohhh, thanks so much :)

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

81 = 3^4 81^(1 / 4) = (3^4)^(1 / 4) = 3^(4 / 4) = 3

OpenStudy (mathstudent55):

Think of "evaluate" as "eVALUate" or finding the value.